OpenStudy (anonymous):

ok so what would N+ 1 over 8 equals 2 over 3 how would i solve n?

OpenStudy (anonymous):

PLEASE HELP!!!!!!

OpenStudy (saifoo.khan):

\[\frac{N+1}{8} = \frac{2}{3}\]

OpenStudy (saifoo.khan):

That?

OpenStudy (anonymous):

no. N is seperate from the 1 over 8 :|

OpenStudy (saifoo.khan):

\[N + \frac18 = \frac23\]

OpenStudy (anonymous):

yes exactly

OpenStudy (saifoo.khan):

\[N = \frac23 -\frac18\]

OpenStudy (anonymous):

?

OpenStudy (saifoo.khan):

Now take LCM, LCM = 24

OpenStudy (saifoo.khan):

??

OpenStudy (anonymous):

o.e uhmm... the lowest common multiple ok well m studying for finals and i suck at math :| and i dont understand wtf u just did :|

OpenStudy (saifoo.khan):

LMAO. i just shifted the 1/8 fraction to the right side. and as we are going to shift it, the sign's gonna change. :p

OpenStudy (anonymous):

hey hey hey ok i suck :P so yeah but thanks :P

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

the answer i would end up getting would be\[13 \over 24\]

OpenStudy (anonymous):

so blechh :P

OpenStudy (anonymous):

aha u fail ;)

OpenStudy (saifoo.khan):